# A Generalized Alternating Anderson Acceleration Method

### v1

@@ 3 Convergence analysis
$\|\mathbf{R}\mathbf{r}_{k}\|$ $=\min_{p\in\mathbb{P}_{m_{k}-1}}\|\left(\mathbf{I}-p(\mathbf{W\Lambda}\mathbf{W}^{-1})\mathbf{W}(\mathbf{I}-\mathbf{\Lambda})\mathbf{\Lambda}^{-m_{k}}\mathbf{W}^{-1}\right)\mathbf{r}_{k}\|$
$\leq\|\mathbf{W}\|\|\mathbf{W}^{-1}\|\min_{p\in\mathbb{P}_{m_{k}-1}}\|\left(\mathbf{I}-p(\mathbf{\Lambda})(\mathbf{I}-\mathbf{\Lambda})\mathbf{\Lambda}^{-m_{k}}\right)\mathbf{r}_{k}\|$
$\leq\kappa_{2}(\mathbf{W})\min_{p\in P_{m_{k}-1}}\max_{t\in[a,b]}|1-p(t)t^{-m_{k}}(1-t)|\|\mathbf{r}_{k}\|$
$\leq\kappa_{2}(\mathbf{W})\min_{p\in P_{m_{k}-1}}\max_{t\in[1/b,1/a]}|1-p(1/t)t^{m_{k}-1}(t-1)|\|\mathbf{r}_{k}\|$
$\leq\kappa_{2}(\mathbf{W})\min_{p\in P_{m_{k}-1}}\max_{t\in[1/b,1/a]}|1-p(t)(t-1)|\|\mathbf{r}_{k}\|$
@@
where the last inequality can be found in the proof of [40, Theorem 3.1]. Because $m\leq k$, $m_{k}=\min\{m,k\}=m$ and we have the desired result (16). Recall that $\mathbf{x}_{k}$ is obtained by applying $k$ times fixed-point iteration to $\mathbf{x}_{0}$. Thus, $\mathbf{r}_{k}=\mathbf{M}^{k}\mathbf{r}_{0}$. It follows that (17) holds. ∎
@@
Note that the $s$ iterates $\{\mathbf{x}_{jp-d}\}_{d=0}^{s-1}$ of aAA(mm)[ss]–FP[$t$] are obtained by AA steps. From [36, Theorem 2.1], we have
@@
Applying Theorem 3 to the (jps+1)(jp-s+1)th iteration, we have
$\|\mathbf{r}_{jp-s+1}\|\leq C(a,b,m)\kappa_{2}(\mathbf{W})\|\mathbf{M}\|\|\mathbf{r}_{jp-s}\|\leq C(a,b,m)\kappa_{2}(\mathbf{W})\|\mathbf{M}\|^{p-s+1}\|\mathbf{r}_{(j-1)p}\|.$
@@
$\|\mathbf{r}_{jp}\|\leq C(a,b,m)\kappa_{2}(\mathbf{W})\|\mathbf{M}\|^{p}\|\mathbf{r}_{(j-1)p}\|\leq(C(a,b,m)\kappa_{2}(\mathbf{W}))^{j}\|\mathbf{M}\|^{jp}\|\mathbf{r}_{0}\|,$
@@
$\min_{p\in P_{k},p(1)=1}\max_{t\in[\alpha,\beta]}|p_{s}(t)|=\frac{1}{|T_{k}(1+2\frac{1-\beta}{\beta-\alpha})|}\leq\epsilon(a,b,k),$
@@
$2\kappa_{2}(\mathbf{W})\left(\frac{\sqrt{\frac{a(1-b)}{b(1-a)}}-1}{\sqrt{\frac{a(1-b)}{b(1-a)}}+1}\right)^{m}b^{t+1}<1,$ (20)
@@ 4 Numerical results
In what follows, the notation $\dagger$ in the tables means that the linear solver did not converge, while the notation $\ddagger$ means that Matlab failed to solve the least-squares problem (4) because the matrix was found to be rank-deficient. Designing efficient solvers for the least-squares problem is left for future work.
@@ 4.2 Application of aAA–FP to nonsymmetric cases
In this subsection, we apply aAA(mm)[ss]–FP[$t$] to solve the nonsymmetric systems $\mathbf{A}\mathbf{x}=\mathbf{b}$ arising from finite elements discretizations. The problems we consider here are from the Matrix Market repository, which is available at https://math.nist.gov/MatrixMarket/. Specifically, we consider the problems: fidap008, fidap029, and fidapm37. We apply AA($m$) and aAA(mm)[ss]–FP[$t$] for different values of $m$, $s$, and $t$. We also consider standard GMRES as a solver. The fixed-point iteration for AA($m$) and aAA(mm)[ss]–FP[$t$] is Jacobi, which is divergent in our cases. For a fair comparison, we run GMRES for the solution of $\mathbf{D}^{-1}\mathbf{Ax}=\mathbf{D}^{-1}\mathbf{b}$, where $\mathbf{D}$ is the diagonal of $\mathbf{A}$, as done in [31]. We allow a maximum of 10,000 iterations and run the solvers until a reduction of $10^{-8}$ in the relative residual is achieved. In Tables 24, we report the number of iterations and the CPU times required for solving the problems considered here; further in Table 2 we report the dimension of each system to be solved together with the condition number of $\mathbf{A}$ and $\mathbf{D}^{-1}\mathbf{A}$.
Table 2: Degrees of freedom $n$, the condition number of $\mathbf{A}$ and of $\mathbf{D}^{-1}\mathbf{A}$, denoted by $\kappa$, number of iterations it, and CPU time for AA(5), aAA(5)[1]–FP[5], and GMRES.
AA(5) aAA(5)[1]–FP[5] GMRES
problem $n$ $\kappa(\mathbf{A})$ $\kappa(\mathbf{D}^{-1}\mathbf{A})$ it CPU it CPU it CPU
fidap008 3096 4.48e+10 7.27e+08 56 0.034 31 0.015 2182 22.9
fidap029 2870 7.52e+02 5.83e+$0$ 17 0.007 19 0.008 22 0.044
fidapm37 9152 8.59e+09 2.84e+07 $\dagger$ $\dagger$ 49 0.16 2375 170
Table 3: Number of iterations it and CPU time for AA($\infty$), AA(10), aAA(10)[1]–FP[4], aAA(10)[1]–FP[6], and aAA(10)[5]–FP[10].
AA($\infty$) AA(10) aAA(10)[1]–FP[4] aAA(10)[1]–FP[6] aAA(10)[5]–FP[10]
@@
fidap008 25 0.029 32 0.034 31 0.022 29 0.019 27 0.023
fidap029 13 0.006 14 0.006 16 0.006 15 0.005 15 0.014
fidapm37 38 0.210 114 0.490 46 0.152 50 0.162 193 0.760
Table 4: Number of iterations it and CPU time for aAA(5)[3]–FP[5], aAA(5)[5]–FP[5], aAA(10)[3]–FP[5], and aAA(10)[5]–FP[5].
aAA(5)[3]–FP[5] aAA(5)[5]–FP[5] aAA(10)[3]–FP[5] aAA(10)[5]–FP[5]
problem it CPU it CPU it CPU it CPU
fidap008 31 0.015 30 0.019 31 0.018 27 0.021
fidap029 17 0.007 17 0.006 15 0.006 17 0.008
fidapm37 71 0.260 57 0.180 47 0.180 47 0.190
From Tables 24, we can observe that aAA(mm)[ss]–FP[$t$] is able to outperform AA($m$), GMRES($m$), and GMRES. In fact, restarted GMRES with restart $m=100$ is not able to converge for the problems fidap008 and fidapm37, and it takes 22 iterations with 0.046 seconds for fidap029. One has to employ full GMRES in order to obtain convergence. On the other hand, AA converges to the prescribed tolerance for all the problems considered, except for fidapm37 when employing a window size $m=5$; this exception arises because the matrix involved in the least-squares problem (4) was found to be rank-deficient. Further, aAA(mm)[ss]–FP[$t$] solves all the problems for all the combinations of the parameters considered here. Finally, comparing the CPU times, we observe that AA($m$) and aAA(mm)[ss]–FP[$t$] are orders of magnitude faster than full GMRES, with aAA(mm)[ss]–FP[$t$] being able to converge in about half of the times of AA($m$) for the problems fidap008 and fidapm37, while the CPU times of the two methods are comparable for the problem fidap029. Although AA($\infty$) is generally not preferred in practice, we report its performance in Table 4 as a reference. We see that (mm)[ss]–FP[$t$] is competitive as AA($\infty$). These tests considered here suggest that aAA(mm)[ss]–FP[$t$] is very promising for solving hard nonsymmetric problems.
@@ 4.3 Application of aAA–FP to the incompressible Navier–Stokes equations
We now test how AA($m$) and aAA(mm)[ss]–FP[$t$] can accelerate the convergence of the Picard iteration. We employ a level of refinement $l=6$, represented by a uniform grid of size $129\times 129$ for the velocity and of size $65\times 65$ for the pressure, while letting varying the viscosity $\nu$. We run AA($m$), aAA(mm)[ss]–FP[$t$], and the Picard iterations until a reduction of $10^{-10}$ in the residual of the fixed point iteration is obtained. For the results presented here, each Picard linearization of the Navier–Stokes equations (that is, the linear system in (21)) is solved exactly, by employing Matlab’s backslash. We report in Figure 7 the results with $\nu=\frac{1}{500}$, and in Figure 8 the results with $\nu=\frac{1}{2500}$, for $m=1$, 5, and 10. From Figures 78 we can observe that the number of iterations required by aAA(mm)[ss]–FP[$t$] is comparable with the number of iterations required by AA($m$) for reaching the prescribed reduction of the nonlinear residual.
@@ 4.4 Application of aAA–FP to nonlinear optimization
ADMM can be reformulated as a fixed-point iteration of the form in (1) by setting $\mathbf{x}_{k}=[\mathbf{y}^{(k)},\widehat{\bm{\lambda}}^{(k)}]$, solving the first step in (22), and then considering $q(\cdot)$ as the result obtained by the second and third steps in (22).
@@ 4.4.1 Total variation
$\min_{\mathbf{x},\mathbf{y}}\dfrac{1}{2}\|\widehat{\mathbf{x}}-\mathbf{y}\|^{2}+\beta\|\mathbf{y}\|^{2}_{1},$
@@
Following the work in [38], we set $n=1000$ and take the vector $\widehat{\mathbf{x}}$ to be randomly generated from the standard normal distribution. Further, we set the smoothing parameter $\beta=0.001\|\widehat{\mathbf{x}}\|_{\infty}$ and the penalty parameter $\mu=10$.
@@
Table 6: Total variation. Number of iterations and CPU time for ADMM, AA, and aAA–FP methods.
@@ 4.4.2 Lasso problem
$\min_{\mathbf{x},\mathbf{y}}\dfrac{1}{2}\|\mathbf{C}\mathbf{x}-\widehat{\mathbf{y}}\|^{2}+\beta\|\mathbf{y}\|_{1},$
@@
Following [38], we set $n_{1}=150$ and $n_{2}=300$, and take matrix $\mathbf{C}$ to be a randomly generated sparse matrix with density 0.01 and the vector $\widehat{\mathbf{x}}$ randomly generated, both sampled from the normal distribution. Further, we set the regularization parameter $\beta=1$ and the penalty parameter $\mu=10$. As above, we allow for 1000 iterations and set as stopping criterion a reduction of $10^{-12}$ in the residual. We report the number of iterations and CPU time in Table 7.
Table 7: Lasso problem. Number of iterations and CPU time for ADMM, AA, and aAA–FP methods.
@@
As we observed for the previous problem, alternating between AA and the fixed-point iteration can result in faster convergence of the method. In fact, aAA(mm)[ss]–FP[$t$] can be between 2 and 3 times faster than AA($m$), compare, for example, aAA(1)[3]–FP[3] and AA(1) in Table 7. More generally, by carefully selecting the parameters, one can obtain a solver that is comparable to AA($\infty$) in terms of both the number of iterations and CPU time; see, e.g., aAA(8)[10]–FP[3] in Table 7. Finally, aAA(mm)[ss]–FP[$t$] can significantly speed up the performance of ADMM; for example, aAA(88)[1010]–FP[33] yields a solver that is over 4 times faster than ADMM in CPU time and requires only one-sixth the number of iterations.

### v2

@@ 3 Convergence analysis for the linear case
$\|\mathbf{R}\mathbf{r}_{k}\|$ $=\min_{p_{\ell}\in\mathbb{P}_{m_{k}-1}}\left\|\left(\mathbf{I}-p_{\ell}(\mathbf{W\Lambda}\mathbf{W}^{-1})\mathbf{W}(\mathbf{I}-\mathbf{\Lambda})\mathbf{\Lambda}^{-m_{k}}\mathbf{W}^{-1}\right)\mathbf{r}_{k}\right\|$
$\leq\|\mathbf{W}\|\|\mathbf{W}^{-1}\|\min_{p_{\ell}\in\mathbb{P}_{m_{k}-1}}\left\|\left(\mathbf{I}-p_{s}(\mathbf{\Lambda})(\mathbf{I}-\mathbf{\Lambda})\mathbf{\Lambda}^{-m_{k}}\right)\right\|\|\mathbf{r}_{k}\|$
$\leq\kappa_{2}(\mathbf{W})\min_{p_{\ell}\in P_{m_{k}-1}}\max_{t\in[a,b]}|1-p_{\ell}(t)t^{-m_{k}}(1-t)|\|\mathbf{r}_{k}\|$
$\leq\kappa_{2}(\mathbf{W})\min_{p_{\ell}\in P_{m_{k}-1}}\max_{t\in[1/b,1/a]}|1-p_{\ell}(1/t)t^{m_{k}-1}(t-1)|\|\mathbf{r}_{k}\|$
$\leq\kappa_{2}(\mathbf{W})\min_{p_{\ell}\in P_{m_{k}-1}}\max_{t\in[1/b,1/a]}|1-p_{\ell}(t)(t-1)|\|\mathbf{r}_{k}\|$
@@
where the last inequality can be found in the proof of [42, Theorem 3.1]. Because $m\leq k$, $m_{k}=\min\{m,k\}=m$ and we have the desired result (19). Recall that $\mathbf{x}_{k}$ is obtained by applying $k$ times fixed-point iteration to $\mathbf{x}_{0}$. Thus, $\mathbf{r}_{k}=\mathbf{M}^{k}\mathbf{r}_{0}$. It follows that (20) holds. ∎
@@
Note that the $s$ iterates $\{\mathbf{x}_{jp-d}\}_{d=0}^{s-1}$ of aAA(mm)[ss]–FP[$t$] are obtained by AA steps. From [38, Theorem 2.1], we have
@@
Note that the $t$ iterates $\{\mathbf{x}_{jp-d}\}_{d=s}^{p-1}$ of aAA(mm)[ss]–FP[$t$] are generated by the fixed-point iterations, while $\mathbf{x}_{jp-s+1}$ is produced by the AA($m$) step, which uses the previous $m+1$ initial guesses. Since $m\leq t$ corresponding to the condition $m\leq k$ in Lemma 1 required by Theorem 3, we can apply Theorem 3 to the (jps+1)(jp-s+1)th iteration. Then, we have
$\|\mathbf{r}_{jp-s+1}\|\leq C(a,b,m)\kappa_{2}(\mathbf{W})\|\mathbf{M}\|\|\mathbf{r}_{jp-s}\|\leq C(a,b,m)\kappa_{2}(\mathbf{W})\|\mathbf{M}\|^{p-s+1}\|\mathbf{r}_{(j-1)p}\|,$
where the first inequality follows from (19), and the second inequality uses the standard property of the error estimation for the fixed-point iterations.
@@
$\|\mathbf{r}_{jp}\|\leq C(a,b,m)\kappa_{2}(\mathbf{W})\|\mathbf{M}\|^{p}\|\mathbf{r}_{(j-1)p}\|\leq(C(a,b,m)\kappa_{2}(\mathbf{W}))^{j}\|\mathbf{M}\|^{jp}\|\mathbf{r}_{0}\|,$ (23)
@@
When $m>t$, the results in Theorem 3 no longer apply. Consequently, we cannot derive the results in Theorem 4. In fact, when $m\leq t$, the $\mathbf{x}_{jp-s+1}$ is a linear combination of $m+1$ fixed-point iterates. While, when $m>t$, the $\mathbf{x}_{jp-s+1}$ is a linear combination of $t$ fixed-point iterates and previous m+1tm+1-t iterates generated by AA($m$), so there is no pattern available to estimate $\mathbf{x}_{jp-s+1}$.
@@
$\min_{p_{\ell}\in P_{k},p_{\ell}(1)=1}\max_{t\in[\alpha,\beta]}|p_{\ell}(t)|=\frac{1}{|T_{k}(1+2\frac{1-\beta}{\beta-\alpha})|}\leq\epsilon(a,b,k),$
@@
$2\kappa_{2}(\mathbf{W})\left(\frac{\sqrt{\frac{a(1-b)}{b(1-a)}}-1}{\sqrt{\frac{a(1-b)}{b(1-a)}}+1}\right)^{m}\|\mathbf{M}\|^{t+s}<1,$ (24)
the subsequence $\{\mathbf{x}_{jp}\}$ generated by aAA(mm)[ss]–FP[$t$] converges. Additionally, if $\mathbf{M}$ is symmetric and
$\left(\frac{\sqrt{\frac{a(1-b)}{b(1-a)}}-1}{\sqrt{\frac{a(1-b)}{b(1-a)}}+1}\right)^{m}b^{t+s}<1,$ (25)
@@
Proof.
From the first inequality in (23), Lemma 2, and the relation $p=t+s$, we derive the first result. If $\mathbf{M}$ is symmetric, then $\|\mathbf{M}\|=b$ and $\kappa_{2}(\mathbf{W})=1$, leading directly to the second result. ∎
@@
From the previous analysis of the linear case, we identify the crucial elements as the properties of the iteration matrix and the use of polynomial approximation techniques. These arguments, however, do not carry over directly to the nonlinear setting.
The main contributions of Theorems 3 and 4 establish general convergence results that do not require contractiveness, and Corollary 1 provides a new sufficient condition for convergence in the non‑contractive case. We believe that these convergence results enrich the current understanding of AA‑type methods and offer guidance for their nonlinear extensions. When the iterates are sufficiently close to the exact solution, the nonlinear iteration can be locally characterized by its linearization, and analyzing this connection will be an important direction for our future work.
@@ 4.2 Application of aAA–FP to nonsymmetric cases
In this subsection, we apply aAA(mm)[ss]–FP[$t$] to solve the nonsymmetric systems $\mathbf{A}\mathbf{x}=\mathbf{b}$ arising from finite elements discretizations. The problems we consider here are from the Matrix Market repository, which is available at https://math.nist.gov/MatrixMarket/. Specifically, we consider the problems: fidap008, fidap029, and fidapm37. We apply AA($m$) and aAA(mm)[ss]–FP[$t$] for different values of $m$, $s$, and $t$. We also consider standard GMRES as a solver. The fixed-point iteration for AA($m$) and aAA(mm)[ss]–FP[$t$] is either weighted Jacobi with weight $\omega=0.5$ or Gauss–Seidel. For a fair comparison, we run GMRES for the solution of $\mathbf{P}^{-1}\mathbf{Ax}=\mathbf{P}^{-1}\mathbf{b}$, where $\mathbf{P}$ is either the diagonal of $\mathbf{A}$ for Jacobi, as done in [34], or its lower-triangular part. We allow a maximum of 5000 iterations and run the solvers until a reduction of $10^{-8}$ in the relative residual is achieved, that is, starting from an initial residual $\mathbf{r}_{0}$, we require that $\|\mathbf{r}_{k}\|/\|\mathbf{r}_{0}\|<10^{-8}$. For all the tests here, we use the vector of all ones as the initial guess. In Tables 2, 3, and 4, we report the number of iterations and the CPU times required for solving the problems considered here; further in Table 2 we report the dimension of each system to be solved together with the condition number of $\mathbf{A}$ and $\mathbf{P}^{-1}\mathbf{A}$ for Jacobi. We note that the Jacobi iteration, viewed as a fixed‑point method, is divergent in all cases, and that the Gauss–Seidel method, also viewed as a fixed‑point iteration, fails to meet the stopping criterion within 5000 iterations in every case.
Table 2: Degrees of freedom $n$, the condition number of $\mathbf{A}$ and of $\mathbf{P}^{-1}\mathbf{A}$, denoted by $\kappa$, number of iterations it, and CPU time for AA(100), aAA(100)[10]–FP[5], and GMRES(100), when applying Jacobi using all ones as the initial guess. The notation $\dagger$ indicates that the solver did not achieve the desired stopping criterion within the 5000 iterations.
AA(100) aAA(100)[10]–FP[5] GMRES(100)
problem $n$ $\kappa(\mathbf{A})$ $\kappa(\mathbf{P}^{-1}\mathbf{A})$ it CPU it CPU it CPU
[…]
fidap029 2870 7.52e+02 5.83e+$0$ 23 0.02 23 0.014 22 0.06
[…]
Table 3: Number of iterations it and CPU time for AA($\infty$), AA(60), aAA(60)[8]–FP[4], aAA(60)[10]–FP[5], and aAA($\infty$)[10]–FP[5], when applying Jacobi using all ones as the initial guess. The notation $\dagger$ indicates that the solver did not achieve the desired stopping criterion within the 5000 iterations. The notation $\ddagger$ means that Matlab reached the maximum allowed CPU time for the run.
AA($\infty$) AA(60) aAA(60)[8]–FP[4] aAA(60)[10]–FP[5] aAA($\infty$)[10]–FP[5]
@@
[…]
fidap029 23 0.022 23 0.02 23 0.015 23 0.015 23 0.016
[…]
Table 4: Number of iterations it and CPU time for AA($\infty$), AA(60), aAA(60)[10]–FP[5], aAA($\infty$)[10]–FP[5], GMRES(60), and GMRES($\infty$), when applying Gauss–Seidel using all ones as the initial guess. The notation $\dagger$ indicates that the solver did not achieve the desired stopping criterion within the 5000 iterations.
AA($\infty$) AA(60) aAA(60)[10]–FP[5] aAA($\infty$)[10]–FP[5] GMRES(60) GMRES($\infty$)
problem it CPU it CPU it CPU it CPU it CPU it CPU
[…]
fidap029 18 0.012 18 0.012 18 0.019 18 0.011 19 0.38 18 0.012
[…]
From Tables 2, 3 and 4, we can observe that aAA(mm)[ss]–FP[$t$] is able to outperform AA($m$) and GMRES($m$) for problem fidap029. In fact, restarted GMRES($m$) with restart $m=100$ takes 22 iterations with 0.06 seconds for this problem. On the other hand, aAA(100)[10]–FP[5] requires 0.014 seconds for reaching convergence, and is about two times faster than AA(100). For problems fidap008 and fidapm37, the solvers are not able to converge to the prescribed accuracy within 5000 iterations (reach $10^{-5}$), and one has to employ full GMRES or window size $m=\infty$ within Anderson approach. This is true also when applying Gauss–Seidel as a fixed-point iteration. We believe that this is the case because the preconditioners employed are not well suited for these two problems. We would like to mention that, for Jacobi, our results differ from the ones presented in [34]; this is because the method presented in [34] includes a relaxation parameter in the Anderson acceleration, which we do not explore.
For these tests, the results suggest that effective preconditioner design, when combined with AA-type methods, is necessary to achieve fast convergence, which we leave for future work. Conversely, AA-type methods may be better suited for nonlinear problems, as demonstrated by our simulations in later sections.
@@ 4.3 Application of aAA–FP to the incompressible Navier–Stokes equations
We now test how AA($m$) and aAA(mm)[ss]–FP[$t$] can accelerate the convergence of the Picard iteration. We employ a level of refinement $l=6$, represented by a uniform grid of size $129\times 129$ for the velocity and of size $65\times 65$ for the pressure, while letting varying the viscosity $\nu$. We run AA($m$), aAA(mm)[ss]–FP[$t$], and the Picard iterations until a reduction of $10^{-10}$ in the residual of the fixed point iteration is obtained. For the results presented here, each Picard linearization of the Navier–Stokes equations (that is, the linear system in (26)) is solved exactly, by employing Matlab’s backslash. We employ the zero vector as the initial guess for the tests in this subsection. We report in Figure 7 the results with $\nu=\frac{1}{500}$, and in Figure 8 the results with $\nu=\frac{1}{2500}$, for $m=1$, 5, and 10. From Figures 78 we can observe that the number of iterations required by aAA(mm)[ss]–FP[$t$] is comparable with the number of iterations required by AA($m$) for reaching the prescribed reduction of the nonlinear residual.
@@ 4.4 Application of aAA–FP to nonlinear optimization
ADMM can be reformulated as a fixed-point iteration of the form in (1) by setting $\mathbf{x}_{k}=[\mathbf{y}^{(k)},\bar{\bm{\lambda}}^{(k)}]$, solving the first step in (27), and then considering $q(\cdot)$ as the result obtained by the second and third steps in (27).
@@ 4.4.1 Total variation
$\min_{\mathbf{x},\mathbf{y}}\dfrac{1}{2}\|\widehat{\mathbf{x}}-\mathbf{x}\|^{2}_{2}+\beta\|\mathbf{y}\|^{2}_{1},$
@@
Following the work in [40], we set $n=1000$ and take the vector $\widehat{\mathbf{x}}$ to be randomly generated from the standard normal distribution. Further, we set the smoothing parameter $\beta=0.001\|\widehat{\mathbf{x}}\|_{\infty}$ and the penalty parameter $\mu=10$. For this test, we employ the zero vector as the initial guess.
@@
Table 6: Total variation. Number of iterations and CPU time for ADMM, AA, and aAA–FP methods. The notation $\dagger$ indicates that the solver did not achieve the desired stopping criterion within 1000 iterations.
@@ 4.4.2 Lasso problem
$\min_{\mathbf{x},\mathbf{y}}\dfrac{1}{2}\|\mathbf{C}\mathbf{x}-\widehat{\mathbf{x}}\|^{2}_{2}+\beta\|\mathbf{y}\|_{1},$
@@
Following [40], we set $n_{1}=150$ and $n_{2}=300$, and take matrix $\mathbf{C}$ to be a randomly generated sparse matrix with density 0.01 and the vector $\widehat{\mathbf{x}}$ randomly generated, both sampled from the normal distribution. Further, we set the regularization parameter $\beta=1$ and the penalty parameter $\mu=10$. As above, we allow for 1000 iterations and set as stopping criterion a reduction of $10^{-12}$ in the residual.
We first consider using the zero vector as the initial guess. We report the number of iterations and CPU time for ADMM, AA, and aAA(mm)[ss]–FP[$t$] methods in Table 7. As we observed for the previous problem, alternating between AA and the fixed-point iteration can result in faster convergence of the method. In Table 7, we observed that aAA(mm)[ss]–FP[$t$] can be between 2 and 3 times faster than AA($m$), compare, for example, aAA(1)[3]–FP[3] and AA(1). More generally, by carefully selecting the parameters, one can obtain a solver that is comparable to AA($\infty$) in terms of both the number of iterations and CPU time; see, e.g., aAA(8)[10]–FP[3] in Table 7. Finally, aAA(mm)[ss]–FP[$t$] can significantly speed up the performance of ADMM; for example, aAA(88)[1010]–FP[33] yields a solver that is over 4 times faster than ADMM in CPU time and requires only one-sixth the number of iterations.
Motivated by observations in the literature that initial guesses influence the performance of AA, we employ 50 different initial guesses generated by Matlab’s rand function with seed 100,000 for the runs. Table 8 reports the average number of iterations and CPU time. Although the average iteration count is higher than the single-run results of Table 7, the average CPU time is lower than that obtained with a single initial guess. Overall, aAA–FP outperforms AA($m$) and ADMM. The results in Table 8 also suggest that different initial guesses have a significant influence on the performance of ADMM, AA, and aAA–FP methods. We do not further investigate the influence of initial guesses on the performance of aAA–FP, which could be an interesting direction for future work.
From Tables 7 and 8, we notice that for fixed $m$ and $t$, increasing $s$ improves the performance, but once $s$ become sufficiently large, the gains diminish. In practice, it is advisable not to choose $s$ to be too large.
Table 7: Lasso problem. Number of iterations and CPU time for ADMM, AA, and aAA–FP methods using the zero vector as the initial guess.
@@
it CPU it CPU it CPU it CPU it CPU
57 0.14 93 0.12 62 0.08 59 0.08 104 0.21
Table 8: Lasso problem. The average of number of iterations and the average of CPU time for ADMM, AA, and aAA–FP methods using 50 random initial guesses generated by Matlab’s rand function with seed 100,000.
ADMM AA(1) AA(3) AA(8) AA($\infty$)
it CPU it CPU it CPU it CPU it CPU
514 0.71 253 0.35 120 0.17 78 0.13 87 0.18
it CPU it CPU it CPU it CPU it CPU
86 0.10 121 0.14 96 0.12 82 0.11 105 0.17
it CPU it CPU it CPU it CPU it CPU
84 0.10 106 0.13 92 0.13 86 0.11 94 0.15
it CPU it CPU it CPU it CPU it CPU
77 0.10 100 0.12 87 0.10 83 0.12 143 0.56
it CPU it CPU it CPU it CPU it CPU
77 0.10 115 0.14 83 0.10 75 0.09 125 0.33
it CPU it CPU it CPU it CPU it CPU
78 0.12 116 0.17 87 0.13 77 0.12 385 4.2